There are x boys and y girls in a class (class A). The girls outnumber the boys by 4. They join up with another class (class B) in which there are twice as many boys as girls. Class B has 3 more boys and 5 less girls than class A. Which systems of equations can be used to solve for x and y?
Answer by Boreal(15235)

x=boys in A
y=girls in A
y-4=x or -x=y-4
Class A has y-4+y students or 2y-4
2x+y is B
x+3 boys and y-5 girls.
But in B x+3=2(y-5)
x=2y-13
x-y-4 rom above.
-x=-y+4 add them
0=y-9
y=9
x=5
Class A has 5 boys and 9 girls
Class B has 8 boys and 4 girls, which is twice as many.